OpenStudy (anonymous):

Ticket for a high school football game were $8 for adults and $5 for students. A total of 1970 tickets were sold. The total sold ticket sales were $12,220. How many adult and student tickets were sold?

hartnn (hartnn):

let there be 'a' adults and 's' students total tickets = tickets for adults and students = a+s = 1970 got this part ?

OpenStudy (anonymous):

yes

hartnn (hartnn):

do you know how to get other equation ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

wait would it be

OpenStudy (anonymous):

8a+5s=12,220

hartnn (hartnn):

yes! absolutely correct :)

hartnn (hartnn):

now we have 2 equations and 2 unknowns can you try to find a and s from these 2 equations ?

hartnn (hartnn):

i'll get you started a+s = 1970 so, a= 1970 - s right ? plug this in 2nd equation what do u get ?

OpenStudy (anonymous):

yea i can do the rest now thanks

OpenStudy (anonymous):

all i needed were how to get the first two equations from the word problem which im really bad at

hartnn (hartnn):

really? good! tell me what values you get so that i can verify them, if you want me to verify :)

OpenStudy (anonymous):

ok let me do it real quick

hartnn (hartnn):

and you're not bad at forming equations, you formed a difficult equation 8a+5s=12,220 correctly! ;)

OpenStudy (anonymous):

lol thanks

hartnn (hartnn):

and i am also quite sure, you'll give me 'a' and 's' values correctly! you seem smart :)

OpenStudy (anonymous):

um would a=2350/3

hartnn (hartnn):

you're quite close actually, how did you get that ?

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

that value of a is not correct...

OpenStudy (anonymous):

yea i figured

OpenStudy (anonymous):

wait let me try it another way

hartnn (hartnn):

take your time :)

OpenStudy (anonymous):

idk i keep getting the same answer 2350/3

hartnn (hartnn):

ok, can you show me what u did ? i'll help you find the error

OpenStudy (anonymous):

8a+5s=12,200 8a -8a 5s=-8a+12,200 /5 /5 /5 s=-8a/5 +2440 a -8a/5 +2440=1970 then just solve for a

hartnn (hartnn):

a+s = 1970 right ? so, a +8a/5 +2440=1970 isn't it ?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

no -8a

hartnn (hartnn):

hmm, yeah, its -8a/5 only

OpenStudy (anonymous):

yea a-8a/5+2440=1970

hartnn (hartnn):

3a/8 = 470 .... why go this complicated way? can i show you another way, which does not involve fractions ?

OpenStudy (anonymous):

sure

hartnn (hartnn):

a+s = 1970 so, a= 1970 - s right ?

hartnn (hartnn):

plug this in 2nd equation, 8a+5s = 12220 what do u get ?

OpenStudy (anonymous):

8(1970-s)+5s=12200

OpenStudy (anonymous):

15760-8s+5s=12200

OpenStudy (anonymous):

thats what i did the first time

OpenStudy (anonymous):

and the end result would be 2350/3

hartnn (hartnn):

15760 - 12220 = 3540 :)

OpenStudy (anonymous):

im getting so confused now

hartnn (hartnn):

15760-8s+5s=12200 15760 - 12220 = 8s-5s = 3s 3s = 3540 s=...

OpenStudy (anonymous):

s=-1180

OpenStudy (anonymous):

1180*

OpenStudy (anonymous):

a=790

hartnn (hartnn):

you are \(\huge \color{red}{\checkmark}\)

OpenStudy (anonymous):

thanks

hartnn (hartnn):

welcome ^_^ oh, and since you are new here, \(\Huge \mathcal{\text{Welcome To OpenStudy}\ddot\smile} \) if you have any doubts browsing this site, you can ask me. Have fun learning with Open Study! :)